# On generating mapping class groups by pseudo-Anosov elements

### v2

@@ 2. Preliminaries
From (6), (7) and the assumption that $f^{-l}(b_{k+3},b_{k+4})=(b_{k+3-l},b_{k+4-l})$ for $l=1,2,3$, we obtain the following:
$t_{b_{k+3}}t_{b_{k+4}}^{-1}=t_{b_{k+3}}t_{c_{k+3}}^{-1}(t_{b_{k+4}}t_{c_{k+3}}^{-1})^{-1}\in G^{\prime\prime},$
$t_{b_{k}}t_{b_{k+1}}^{-1},t_{b_{k+1}}t_{b_{k+2}}^{-1},t_{b_{k+2}}t_{b_{k+3}}^{-1}\in G^{\prime\prime}.$
Moreover, by (1), (2), $t_{b_{k+2}}t_{b_{k+3}}^{-1}\in G^{\prime\prime}$ and the assumption that $f^{-2}(a_{k+2},a_{k+3})=(a_{k},a_{k+1})$, we have
$t_{a_{k+2}}t_{a_{k+3}}^{-1}=t_{a_{k+2}}t_{b_{k+2}}^{-1}\cdot t_{b_{k+2}}t_{b_{k+3}}^{-1}\cdot(t_{a_{k+3}}t_{b_{k+3}}^{-1})^{-1}\in G^{\prime\prime},$
$t_{a_{k}}t_{a_{k+1}}\in G^{\prime\prime}.$
By (8), (6) and the assumption that $f^{-2}(c_{k+2},c_{k+3})=(c_{k},c_{k+1})$, the following holds:
$t_{c_{k+2}}t_{c_{k+3}}^{-1}=(t_{b_{k+3}}t_{c_{k+2}}^{-1})^{-1}t_{b_{k+3}}t_{c_{k+3}}^{-1}\in G^{\prime\prime},$
$t_{c_{k}}t_{c_{k+1}}^{-1}\in G^{\prime\prime}.$
Summarizing the above, the elements $t_{a_{k}}t_{a_{k+1}}^{-1},t_{b_{k}}t_{b_{k+1}}^{-1},t_{c_{k}}t_{c_{k+1}}$ are contained in $G^{\prime\prime}$, and the lemma follows. ∎
@@
Since $a_{1},b_{1},c_{1}$ are disjoint from $b_{3},a_{5},c_{6}$, we see that
@@
$=(t_{b_{3}}t_{a_{5}}^{-1}t_{c_{6}})\cdot(t_{\rho_{n}(b_{3})}t_{\rho_{n}(a_{5})}^{-1}t_{\rho_{n}(c_{6})})^{-1}\cdot\rho_{n}$
$=(t_{b_{3}}t_{a_{5}}^{-1}t_{c_{6}})(t_{b_{4}}t_{a_{6}}^{-1}t_{c_{7}})^{-1}\rho_{n}.$
and therefore, by $\rho_{n},(t_{b_{3}}t_{a_{5}}^{-1}t_{c_{6}})\rho_{n}(t_{b_{3}}t_{a_{5}}^{-1}t_{c_{6}})^{-1}\in G_{1}$, $G_{1}$ contains the element $(t_{b_{3}}t_{a_{5}}^{-1}t_{c_{6}})(t_{b_{4}}t_{a_{6}}^{-1}t_{c_{7}})^{-1}$. From Lemma 4.7, we have $t_{a_{3}}t_{a_{4}}^{-1},t_{b_{3}}t_{b_{4}}^{-1},t_{c_{3}}t_{c_{4}}^{-1}\in G_{1}$ (by letting $k=3$ and $f=\rho_{n}$). Moreover, from Lemma 4.6, we see that the elements $t_{a_{3}},t_{b_{3}},t_{c_{3}}$ are in $G_{1}$ (by letting $k=3$ and $f=\rho_{n}$). Since it is easy to check that $\rho_{n}(a_{i},b_{i})=(a_{i+1},b_{i+1})$ for $i=3,4,\ldots,g-1$ and $\rho_{n}(a_{g},b_{g})=(a_{1},b_{1})$, by $t_{a_{3}},t_{b_{3}},\rho_{n}^{g-2}\in G_{1}$, we have
$t_{a_{1}}\in G_{1},$
$t_{b_{1}}\in G_{1}.$
Hence, we obtain $rt_{c_{1}}=rt_{c_{1}}t_{b_{1}}^{-n}t_{a_{1}}\cdot t_{a_{1}}^{-1}t_{b_{1}}^{n}=\rho_{n}t_{a_{1}}^{-1}t_{b_{1}}^{n}\in G_{1}$. By $(rt_{c_{1}})^{-2}(c_{3})=c_{1}$, we have
$t_{c_{1}}$ $=t_{(rt_{c_{1}})^{-2}(c_{3})}=(rt_{c_{1}})^{-2}t_{c_{3}}(rt_{c_{1}})^{2}\in G_{1},$
$r$ $=rt_{c_{1}}\cdot t_{c_{1}}^{-1}\in G_{1}.$
Since $r^{i}(a_{1},b_{1},c_{1})=(a_{i+1},b_{i+1},c_{i+1})$ for $i=0,1,2,\ldots,g-1$, we have
$t_{a_{i+1}},t_{b_{i+1}},t_{c_{i+1}}\in G_{1}$
for $i=0,1,2,\ldots,g-1$. Therefore, we see that $G_{1}=\mathcal{M}_{g}$, which proves the theorem. ∎
@@
Since $a_{1},b_{1},c_{1}$ are disjoint from $a_{2},c_{2},a_{3},b_{3}$, we see that $\rho_{n}(a_{2},c_{2},a_{3},b_{3})=rt_{c_{1}}t_{b_{1}}^{-n}t_{a_{1}}(a_{2},a_{3},b_{3},c_{2})=r(a_{2},a_{3},b_{3},c_{2})$, and hence
$\rho_{n}(a_{2})$ $=a_{3},$
$\rho_{n}(a_{3},b_{3},c_{2})$ =$=\begin{cases}(a_{4},b_{4},c_{3})&(g\geq 4),\\ (a_{1},b_{1},c_{3})&(g=3).\end{cases}$
@@
$(t_{a_{3}}t_{b_{3}}t_{c_{2}})\rho_{n}(t_{a_{3}}t_{b_{3}}t_{c_{2}})^{-1}$ $=(t_{a_{3}}t_{b_{3}}t_{c_{2}})\{\rho_{n}(t_{a_{3}}t_{b_{3}}t_{c_{2}})^{-1}\rho_{n}^{-1}\}\rho_{n},$
from the above arguments and the relation $t_{f(c)}=ft_{c}f^{-1}$, the following holds:
$G_{2}\ni t_{a_{2}}\rho_{n}t_{a_{2}}^{-1}=t_{a_{2}}t_{a_{3}}^{-1}\rho_{n},$
@@
$t_{a_{3}}t_{b_{3}}t_{c_{2}}\cdot t_{c_{3}}^{-1}t_{b_{4}}^{-1}t_{a_{4}}^{-1}\in G_{2}\ \mathrm{if}\ g\geq 4,$
$t_{a_{3}}t_{b_{3}}t_{c_{2}}\cdot t_{c_{3}}^{-1}t_{b_{1}}^{-1}t_{a_{1}}^{-1}\in G_{2}\ \mathrm{if}\ g=3.$
Hereafter, we prove that $t_{a_{3}}\in G_{2}$ using these elements and Lemma 4.5. More precisely, we show that
$t_{a_{1}}t_{a_{2}}^{-1},\ t_{a_{1}}t_{a_{3}}^{-1},\ t_{a_{1}}t_{b_{2}}^{-1},\ t_{a_{1}}t_{c_{1}}^{-1},\ t_{a_{2}}t_{c_{2}}^{-1},\ t_{a_{1}}t_{c_{2}}^{-1},\ t_{a_{1}}t_{b_{3}}^{-1}\in G_{2}.$
Suppose that $g\geq 4$. For simplicity of notation, we write $\phi=t_{a_{3}}t_{b_{3}}t_{c_{2}}\cdot t_{c_{3}}^{-1}t_{b_{4}}^{-1}t_{a_{4}}^{-1}(\in G_{2})$.
We first show that $t_{a_{1}}t_{a_{2}}^{-1},t_{a_{1}}t_{a_{3}}^{-1},t_{a_{1}}t_{b_{3}}^{-1}\in G_{2}$. It is easy to check that $\phi(a_{2},a_{3})=(a_{2},b_{3})$, and hence we have $t_{a_{2}}t_{b_{3}}^{-1}\in G_{2}$ by $t_{a_{2}}t_{a_{3}}^{-1},\phi\in G_{2}$. Since $a_{1},b_{1},c_{1}$ are disjoint from $a_{i_{1}}$ and $b_{i_{2}}$ for $i_{1}=2,3,\ldots,g$ and $i_{2}=3,4,\ldots,g$, we see that $\rho_{n}(a_{i},a_{i+1},b_{i+1})=rt_{c_{1}}t_{b_{1}}^{-n}t_{a_{1}}(a_{i},a_{i+1},b_{i+1})=(a_{i+1},a_{i+2},b_{i+2})$ for $i=2,3,\ldots,g-2$ and $\rho_{n}(a_{g},b_{g})=rt_{c_{1}}t_{b_{1}}^{-n}t_{a_{1}}(a_{g},b_{g})=(a_{1},b_{1})$. From the above arguments and $\rho_{n}^{i},t_{a_{2}}t_{a_{3}}^{-1},t_{a_{2}}t_{b_{3}}^{-1}\in G_{2}$ for $i=1,2,\ldots,g-2$, we have
(10) $t_{a_{2+i}}t_{a_{3+i}}^{-1}\in G_{2}\ \ (0\leq i\leq g-3)$ $\mathrm{and}$ $t_{a_{g}}t_{a_{1}}^{-1}\in G_{2}\ (i=g-2),$
(11) $t_{a_{g}}t_{b_{1}}^{-1}\in G_{2}\ (i=g-2).$
Using these, we have
$t_{a_{2}}t_{a_{1}}^{-1}=t_{a_{2}}t_{a_{3}}^{-1}\cdot t_{a_{3}}t_{a_{4}}^{-1}\cdots t_{a_{g-1}}t_{a_{g}}^{-1}\cdot t_{a_{g}}t_{a_{1}}^{-1}\in G_{2}.$
Since $a_{1}$ (resp. $a_{g}$) is disjoint from $a_{2}$ (resp. $a_{1}$ and $b_{1}$), the following holds:
@@
Moreover, since $\rho_{n}^{-g+2}(a_{1},b_{1})=(a_{3},b_{3})$, we have
(15) $t_{a_{3}}t_{b_{3}}^{-1}=t_{\rho_{n}^{-g+2}(a_{1})}t_{\rho_{n}^{-g+2}(b_{1})}^{-1}=\rho_{n}^{-g+2}t_{a_{1}}t_{b_{1}}^{-1}\rho_{n}^{g-2}\in G_{2}.$
Therefore, we see that
@@
From (12), (13) and (16), we have $t_{a_{1}}t_{a_{2}}^{-1},t_{a_{1}}t_{a_{3}}^{-1},t_{a_{1}}t_{b_{3}}^{-1}\in G_{2}$.
Next, we show that $t_{a_{1}}t_{b_{2}}^{-1},t_{a_{1}}t_{c_{1}}^{-1},t_{a_{2}}t_{c_{2}}^{-1},t_{a_{1}}t_{c_{2}}^{-1}\in G_{2}$. By (9) and (15), we have $t_{a_{2}}t_{b_{3}}^{-1}=t_{a_{2}}t_{a_{3}}^{-1}\cdot t_{a_{3}}t_{b_{3}}^{-1}\in G_{2}$, and hence, by $\rho_{n}(a_{2},b_{3})=(a_{3},b_{4})$ and $t_{a_{2}}t_{b_{3}}^{-1},\rho_{n}\in G_{2}$, we have $t_{a_{3}}t_{b_{4}}^{-1}\in G_{2}$. Combining this and (13), the element $t_{a_{1}}t_{b_{4}}^{-1}=t_{a_{1}}t_{a_{3}}^{-1}\cdot t_{a_{3}}t_{b_{4}}^{-1}$ is in $G_{2}$. Since it is easily seen that $\phi^{-1}(a_{1},b_{4})=(a_{1},c_{3})$, we have $t_{a_{1}}t_{c_{3}}^{-1}\in G_{2}$ by $t_{a_{1}}t_{b_{4}}^{-1},\phi^{-1}\in G_{2}$. Using this, (12) and (13), we have
$t_{a_{2}}t_{c_{3}}^{-1}=(t_{a_{1}}t_{a_{2}}^{-1})^{-1}t_{a_{1}}t_{c_{3}}^{-1}\in G_{2},$
$t_{a_{3}}t_{c_{3}}^{-1}=(t_{a_{1}}t_{a_{3}}^{-1})^{-1}t_{a_{1}}t_{c_{3}}^{-1}\in G_{2}.$
Here, we set $\sigma_{n}=\rho_{n}\cdot t_{a_{1}}^{-1}t_{b_{1}}\cdot(t_{a_{2}}^{-1}t_{a_{1}}\cdot t_{a_{1}}^{-1}t_{b_{1}})^{n-1}$. Then, by the disjointness of $a_{2}$ and $b_{1}$, (12) and (14), we have $G_{2}\ni\sigma_{n}=rt_{c_{1}}t_{a_{2}}^{-n+1}$. Since it is obvious that $\sigma_{n}^{-2}(a_{3},c_{3})=(a_{1},c_{1})$, $\sigma_{n}^{-1}(a_{3},c_{3})=(a_{2},c_{2})$ and $\sigma_{n}^{-2}(a_{2},c_{3})=(a_{1},c_{2})$, by $t_{a_{3}}t_{c_{3}}^{-1},t_{a_{2}}t_{c_{3}}^{-1},\sigma_{n}\in G_{2}$, we see that the following holds:
@@
Moreover, since it is immediate that $\sigma_{n}(a_{1},b_{1})=(a_{2},t_{c_{2}}(b_{2}))$, by (14) and $\sigma_{n}\in G_{2}$ we have $t_{a_{2}}^{-1}t_{t_{c_{2}}(b_{2})}\in G_{2}$. Since $t_{a_{2}}^{-1}t_{t_{c_{2}}(b_{2})}=t_{a_{2}}^{-1}t_{c_{2}}t_{b_{2}}t_{c_{2}}^{-1}$ by the relation $t_{t_{c_{2}}(b_{2})}=t_{c_{2}}t_{b_{2}}t_{c_{2}}^{-1}$, we have $t_{c_{2}}t_{b_{2}}^{-1}\in G_{2}$ by (18) and the disjointness of $a_{2}$ and $c_{2}$. Therefore, by (19) the element $t_{a_{1}}t_{b_{2}}^{-1}=t_{a_{1}}t_{c_{2}}^{-1}\cdot t_{c_{2}}t_{b_{2}}^{-1}$ is contained in $G_{2}$. Summarizing the above arguments, we have $t_{a_{1}}t_{b_{2}}^{-1},t_{a_{1}}t_{c_{1}}^{-1},t_{a_{2}}t_{c_{2}}^{-1},t_{a_{1}}t_{c_{2}}^{-1}\in G_{2}$.
By Lemma 4.5 we see that $t_{a_{3}}\in G_{2}$. Moreover, by (13), (14) and (17) we have
$t_{a_{1}}$ $=t_{a_{1}}t_{a_{3}}^{-1}\cdot t_{a_{3}}\in G_{2},$
$t_{b_{1}}$ $=t_{a_{1}}\cdot t_{a_{1}}^{-1}t_{b_{1}}\in G_{2},$
$t_{c_{1}}$ $=(t_{a_{1}}t_{c_{1}}^{-1})^{-1}\cdot t_{a_{1}}^{-1}\in G_{2},$
and hence $r=\rho_{n}t_{a_{1}}^{-1}t_{b_{1}}^{n}t_{c_{1}}^{-1}\in G_{2}$. Since $r^{i}(a_{1},b_{1},c_{1})=(a_{i+1},b_{i+1},c_{i+1})$ for $i=0,1,2,\ldots,g-1$, we have
$t_{a_{i+1}},t_{b_{i+1}},t_{c_{i+1}}\in G_{2}$
for $i=0,1,2,\ldots,g-1$. Therefore, we see that $G_{2}=\mathcal{M}_{g}$ for $g\geq 4$.
Suppose that $g=3$. For simplicity of notation, we also write $\phi=t_{a_{3}}t_{b_{3}}t_{c_{2}}\cdot t_{c_{3}}^{-1}t_{b_{1}}^{-1}t_{a_{1}}^{-1}(\in G_{2})$. Since it is easy to check that $\rho_{n}(a_{2},a_{3})=(a_{3},a_{1})$, by (9) and $\rho_{n}\in G_{2}$ we have
(20) $t_{a_{3}}t_{a_{1}}^{-1}\in G_{2},$
and hence by (9), $\phi(a_{2},a_{3})=(a_{2},b_{3})$ and $\phi\in G_{2}$, we see that $t_{a_{2}}t_{b_{3}}^{-1}\in G_{2}$. From this, $\rho_{n}(a_{2},b_{3})=(a_{3},b_{1})$ and $\rho_{n}\in G_{2}$, we have
@@
Moreover, the following holds:
(22) $t_{a_{1}}t_{a_{3}}^{-1}=(t_{a_{3}}t_{a_{1}}^{-1})^{-1}\in G_{2},$
(23) $t_{a_{1}}t_{a_{2}}^{-1}=t_{a_{1}}t_{a_{3}}^{-1}\cdot(t_{a_{2}}t_{a_{3}}^{-1})^{-1}\in G_{2},$
(24) $t_{a_{1}}t_{b_{3}}^{-1}=t_{a_{1}}t_{a_{2}}^{-1}\cdot t_{a_{2}}t_{b_{3}}^{-1}\in G_{2}.$
By (9), (20), (21) and the disjointness of $a_{3}$ and $a_{1},b_{1}$, we have
(25) $t_{a_{1}}^{-1}t_{b_{1}}=t_{a_{3}}t_{a_{1}}^{-1}\cdot(t_{a_{3}}t_{b_{1}}^{-1})^{-1}\in G_{2},$
(26) $t_{a_{2}}t_{b_{1}}^{-1}=t_{a_{2}}t_{a_{3}}^{-1}\cdot t_{a_{3}}t_{b_{1}}^{-1}\in G_{2}$
Here, when we set $\sigma_{n}=\rho_{n}\cdot t_{a_{1}}^{-1}t_{b_{1}}(t_{a_{2}}t_{b_{1}}^{-1})^{-n+1}$, we see that $\sigma_{n}=rt_{c_{1}}t_{a_{2}}^{-n+1}\in G_{2}$ by (26) and the disjointness of $a_{2}$ and $b_{1}$. Since it is obvious that $\phi^{-1}(a_{2},b_{1})=(a_{2},c_{3})$, by (26) and $\phi^{-1}\in G_{2}$ we have $t_{a_{2}}t_{c_{3}}^{-1}\in G_{2}$. Moreover, by $\sigma_{n}(a_{2},c_{3})=(a_{3},c_{1}),\sigma_{n}(a_{3},c_{1})=(a_{1},c_{2})$ and $t_{a_{2}}t_{c_{3}}^{-1},\sigma_{n}\in G_{2}$ we see that
$t_{a_{3}}t_{c_{1}}^{-1}\in G_{2},$

### v3

@@ 2. Preliminaries
Here, by (6) and (7), the element $t_{b_{k+3}}t_{b_{k+4}}^{-1}=t_{b_{k+3}}t_{c_{k+3}}^{-1}(t_{b_{k+4}}t_{c_{k+3}}^{-1})^{-1}$ is in $G^{\prime\prime}$. Therefore, by the assumption that $f^{-l}(b_{k+3},b_{k+4})=(b_{k+3-l},b_{k+4-l})$ for $l=1,2,3$, $t_{b_{k}}t_{b_{k+1}}^{-1}$, $t_{b_{k+1}}t_{b_{k+2}}^{-1}$ and $t_{b_{k+2}}t_{b_{k+3}}^{-1}$ are contained in $G^{\prime\prime}$. Moreover, by (1), (2), we obtain that $t_{a_{k+2}}t_{a_{k+3}}^{-1}=t_{a_{k+2}}t_{b_{k+2}}^{-1}\cdot t_{b_{k+2}}t_{b_{k+3}}^{-1}\cdot(t_{a_{k+3}}t_{b_{k+3}}^{-1})^{-1}$ is in $G^{\prime\prime}$. By the assumption $f^{-2}(a_{k+2},a_{k+3})=(a_{k},a_{k+1})$, we obtain that $t_{a_{k}}t_{a_{k+1}}^{-1}\in G^{\prime\prime}$. By (8), (6), the element $t_{c_{k+2}}t_{c_{k+3}}^{-1}=(t_{b_{k+3}}t_{c_{k+2}}^{-1})^{-1}t_{b_{k+3}}t_{c_{k+3}}^{-1}$ is in $G^{\prime\prime}$, and therefore $t_{c_{k}}t_{c_{k+1}}^{-1}$ is in $G^{\prime\prime}$ by the assumption that $f^{-2}(c_{k+2},c_{k+3})=(c_{k},c_{k+1})$. Summarizing the above, the elements $t_{a_{k}}t_{a_{k+1}}^{-1},t_{b_{k}}t_{b_{k+1}}^{-1},t_{c_{k}}t_{c_{k+1}}$ are contained in $G^{\prime\prime}$, and the lemma follows. ∎
@@
It is easy to check that
@@
$=(t_{b_{3}}t_{a_{5}}^{-1}t_{c_{6}})\cdot(t_{b_{4}}t_{a_{6}}^{-1}t_{c_{7}})^{-1}\rho_{n}.$
and therefore, $G_{1}$ contains the element $(t_{b_{3}}t_{a_{5}}^{-1}t_{c_{6}})(t_{b_{4}}t_{a_{6}}^{-1}t_{c_{7}})^{-1}$. From Lemma 4.7, the elements $t_{a_{3}}t_{a_{4}}^{-1}$, $t_{b_{3}}t_{b_{4}}^{-1}$ and $t_{c_{3}}t_{c_{4}}^{-1}$ are in $G_{1}$ (by letting $k=3$ and $f=\rho_{n}$). Moreover, from Lemma 4.6, we see that the elements $t_{a_{3}},t_{b_{3}},t_{c_{3}}$ are in $G_{1}$ (by letting $k=3$ and $f=\rho_{n}$). Since it is easy to check that $\rho_{n}(a_{i},b_{i})=(a_{i+1},b_{i+1})$ for $i=3,4,\ldots,g-1$ and $\rho_{n}(a_{g},b_{g})=(a_{1},b_{1})$, we see that $t_{a_{1}}$ and $t_{b_{1}}$ are contained in $G_{1}$. Hence, the element $rt_{c_{1}}=rt_{c_{1}}t_{b_{1}}^{-n}t_{a_{1}}\cdot t_{a_{1}}^{-1}t_{b_{1}}^{n}=\rho_{n}t_{a_{1}}^{-1}t_{b_{1}}^{n}$ is in $G_{1}$. By $(rt_{c_{1}})^{-2}(c_{3})=c_{1}$, we obtain that $t_{c_{1}}=t_{(rt_{c_{1}})^{-2}(c_{3})}=(rt_{c_{1}})^{-2}t_{c_{3}}(rt_{c_{1}})^{2}$ is in $G_{1}$, and therefore, $r=rt_{c_{1}}\cdot t_{c_{1}}^{-1}$ is in $G_{1}$. Since $r^{i}(a_{1},b_{1},c_{1})=(a_{i+1},b_{i+1},c_{i+1})$ for $i=0,1,2,\ldots,g-1$, the elements $t_{a_{i+1}}$, $t_{b_{i+1}}$ and $t_{c_{i+1}}$ are in $G_{1}$ for $i=0,1,2,\ldots,g-1$. Therefore, we see that $G_{1}=\mathcal{M}_{g}$, which proves the theorem. ∎
@@
By the definition of $\rho_{n}$, we obtain
$\rho_{n}(a_{2})$ $=a_{3},\ \ \mathrm{and}$
$\rho_{n}(a_{3},b_{3},c_{2})$ $=\begin{cases}(a_{4},b_{4},c_{3})&(g\geq 4),\\ (a_{1},b_{1},c_{3})&(g=3).\end{cases}$
@@
$(t_{a_{3}}t_{b_{3}}t_{c_{2}})\rho_{n}(t_{a_{3}}t_{b_{3}}t_{c_{2}})^{-1}$ $=(t_{a_{3}}t_{b_{3}}t_{c_{2}})\cdot\rho_{n}(t_{a_{3}}t_{b_{3}}t_{c_{2}})^{-1}\rho_{n}^{-1}\cdot\rho_{n},$
the following holds:
$G_{2}\ni t_{a_{2}}\rho_{n}t_{a_{2}}^{-1}=t_{a_{2}}t_{a_{3}}^{-1}\rho_{n},\ \ \mathrm{and}$
@@
(10) $t_{a_{3}}t_{b_{3}}t_{c_{2}}\cdot t_{c_{3}}^{-1}t_{b_{4}}^{-1}t_{a_{4}}^{-1}\in G_{2}\ \mathrm{if}\ g\geq 4,\ \ \mathrm{and}$
(11) $t_{a_{3}}t_{b_{3}}t_{c_{2}}\cdot t_{c_{3}}^{-1}t_{b_{1}}^{-1}t_{a_{1}}^{-1}\in G_{2}\ \mathrm{if}\ g=3.$
Hereafter, we prove that $t_{a_{3}}$ is in $G_{2}$ using these elements and Lemma 4.5. More precisely, we show that the following elements are contained in $G_{2}$:
$t_{a_{1}}t_{a_{2}}^{-1},\ t_{a_{1}}t_{a_{3}}^{-1},\ t_{a_{1}}t_{b_{2}}^{-1},\ t_{a_{1}}t_{c_{1}}^{-1},\ t_{a_{2}}t_{c_{2}}^{-1},\ t_{a_{1}}t_{c_{2}}^{-1},\ t_{a_{1}}t_{b_{3}}^{-1}.$
Suppose that $g\geq 4$. For simplicity of notation, we write $\phi=t_{a_{3}}t_{b_{3}}t_{c_{2}}\cdot t_{c_{3}}^{-1}t_{b_{4}}^{-1}t_{a_{4}}^{-1}$.
We first show that $t_{a_{1}}t_{a_{2}}^{-1}$, $t_{a_{1}}t_{a_{3}}^{-1}$ and $t_{a_{1}}t_{b_{3}}^{-1}$ are in $G_{2}$. It is easy to check that $\phi(a_{2},a_{3})=(a_{2},b_{3})$, and hence $t_{a_{2}}t_{b_{3}}^{-1}$ is in $G_{2}$ by (9) and (10). We obtain the following:
(12) $t_{a_{2+i}}t_{a_{3+i}}^{-1}=\rho_{n}^{i}t_{a_{2}}t_{a_{3}}^{-1}\rho_{n}^{-i}\in G_{2}\ \ (0\leq i\leq g-3),$
(13) $t_{a_{g}}t_{a_{1}}^{-1}=\rho_{n}^{g-2}t_{a_{2}}t_{a_{3}}^{-1}\rho_{n}^{-(g-2)}\in G_{2},$
(14) $t_{a_{g}}t_{b_{1}}^{-1}=\rho_{n}^{g-2}t_{a_{2}}t_{b_{3}}^{-1}\rho_{n}^{-(g-2)}\in G_{2}.$
Using these, we see that the element $t_{a_{2}}t_{a_{1}}^{-1}=t_{a_{2}}t_{a_{3}}^{-1}\cdot t_{a_{3}}t_{a_{4}}^{-1}\cdots t_{a_{g-1}}t_{a_{g}}^{-1}\cdot t_{a_{g}}t_{a_{1}}^{-1}$ is in $G_{2}$. Therefore, the following holds:
@@
Moreover, since $\rho_{n}^{-g+2}(a_{1},b_{1})=(a_{3},b_{3})$, we have the following:
(18) $t_{a_{3}}t_{b_{3}}^{-1}\in G_{2},\ \ \mathrm{and}$
@@
From (15), (16) and (19), the elements $t_{a_{1}}t_{a_{2}}^{-1}$, $t_{a_{1}}t_{a_{3}}^{-1}$ and $t_{a_{1}}t_{b_{3}}^{-1}$ are in $G_{2}$.
Next, we show that $t_{a_{1}}t_{b_{2}}^{-1}$, $t_{a_{1}}t_{c_{1}}^{-1}$, $t_{a_{2}}t_{c_{2}}^{-1}$ and $t_{a_{1}}t_{c_{2}}^{-1}$ are in $G_{2}$. By (9) and (18), we have $t_{a_{2}}t_{b_{3}}^{-1}=t_{a_{2}}t_{a_{3}}^{-1}\cdot t_{a_{3}}t_{b_{3}}^{-1}$ is in $G_{2}$. Therefore, by $\rho_{n}(a_{2},b_{3})=(a_{3},b_{4})$, the element $t_{a_{3}}t_{b_{4}}^{-1}$ is contained in $G_{2}$. Combining this and (16), the element $t_{a_{1}}t_{b_{4}}^{-1}=t_{a_{1}}t_{a_{3}}^{-1}\cdot t_{a_{3}}t_{b_{4}}^{-1}$ is in $G_{2}$. Then, since it is easily seen that $\phi^{-1}(a_{1},b_{4})=(a_{1},c_{3})$, we see that $t_{a_{1}}t_{c_{3}}^{-1}$ is in $G_{2}$ by (10). Using this, (15) and (16), the elements $t_{a_{2}}t_{c_{3}}^{-1}=(t_{a_{1}}t_{a_{2}}^{-1})^{-1}t_{a_{1}}t_{c_{3}}^{-1}$ and $t_{a_{3}}t_{c_{3}}^{-1}=(t_{a_{1}}t_{a_{3}}^{-1})^{-1}t_{a_{1}}t_{c_{3}}^{-1}$ are contained in $G_{2}$. Here, we set $\sigma_{n}=\rho_{n}\cdot t_{a_{1}}^{-1}t_{b_{1}}\cdot(t_{a_{2}}^{-1}t_{a_{1}}\cdot t_{a_{1}}^{-1}t_{b_{1}})^{n-1}$, which is in $G_{2}$ by (15) and (17). Then, by the definition of $\rho_{n}$, we have $\sigma_{n}=rt_{c_{1}}t_{a_{2}}^{-n+1}$. This gives $\sigma_{n}^{-2}(a_{3},c_{3})=(a_{1},c_{1})$, $\sigma_{n}^{-1}(a_{3},c_{3})=(a_{2},c_{2})$ and $\sigma_{n}^{-2}(a_{2},c_{3})=(a_{1},c_{2})$. Therefore, the following holds:
@@
Moreover, since it is immediate that $\sigma_{n}(a_{1},b_{1})=(a_{2},t_{c_{2}}(b_{2}))$, the element $t_{a_{2}}^{-1}t_{t_{c_{2}}(b_{2})}$ is in $G_{2}$ by (17). By (21) and $t_{a_{2}}^{-1}t_{t_{c_{2}}(b_{2})}=t_{a_{2}}^{-1}t_{c_{2}}t_{b_{2}}t_{c_{2}}^{-1}$, we see that $t_{c_{2}}t_{b_{2}}^{-1}$ is in $G_{2}$. Therefore, the element $t_{a_{1}}t_{b_{2}}^{-1}=t_{a_{1}}t_{c_{2}}^{-1}\cdot t_{c_{2}}t_{b_{2}}^{-1}$ is contained in $G_{2}$ by (22). Summarizing the above, the elements $t_{a_{1}}t_{b_{2}}^{-1}$, $t_{a_{1}}t_{c_{1}}^{-1}$, $t_{a_{2}}t_{c_{2}}^{-1}$ and $t_{a_{1}}t_{c_{2}}^{-1}$ are in $G_{2}$.
By Lemma 4.5 we see that $t_{a_{3}}$ is in $G_{2}$. Moreover, by (16), (17) and (20), the Dehn twists $t_{a_{1}}=t_{a_{1}}t_{a_{3}}^{-1}\cdot t_{a_{3}}$, $t_{b_{1}}=t_{a_{1}}\cdot t_{a_{1}}^{-1}t_{b_{1}}$ and $t_{c_{1}}=(t_{a_{1}}t_{c_{1}}^{-1})^{-1}\cdot t_{a_{1}}^{-1}$ are contained in $G_{2}$, and hence the rotation $r=\rho_{n}t_{a_{1}}^{-1}t_{b_{1}}^{n}t_{c_{1}}^{-1}$ is in $G_{2}$. Since $r^{i}(a_{1},b_{1},c_{1})=(a_{i+1},b_{i+1},c_{i+1})$ for $i=0,1,2,\ldots,g-1$, we obtain that the Dehn twists $t_{a_{i+1}}$, $t_{b_{i+1}}$ and $t_{c_{i+1}}$ are in $G_{2}$ for $i=0,1,2,\ldots,g-1$. Therefore, we see that $G_{2}=\mathcal{M}_{g}$ for $g\geq 4$.
Suppose that $g=3$. Since it is easy to check that $\rho_{n}(a_{2},a_{3})=(a_{3},a_{1})$, by (9) the following holds:
(23) $t_{a_{3}}t_{a_{1}}^{-1}=t_{a_{1}}^{-1}t_{a_{3}}\in G_{2},\ \ \mathrm{and}$
(24) $t_{a_{1}}t_{a_{2}}^{-1}=(t_{a_{3}}t_{a_{1}}^{-1})^{-1}\cdot(t_{a_{2}}t_{a_{3}}^{-1})^{-1}\in G_{2}.$
Here, for simplicity of notation, we also write $\phi=t_{a_{3}}t_{b_{3}}t_{c_{2}}\cdot t_{c_{3}}^{-1}t_{b_{1}}^{-1}t_{a_{1}}^{-1}$. Then, by (9), (11) and $\phi(a_{2},a_{3})=(a_{2},b_{3})$, we see that $t_{a_{2}}t_{b_{3}}^{-1}$ is in $G_{2}$. From this, (24) and $\rho_{n}(a_{2},b_{3})=(a_{3},b_{1})$, we have the following:
(25) $t_{a_{1}}t_{b_{3}}^{-1}=t_{a_{1}}t_{a_{2}}^{-1}\cdot t_{a_{2}}t_{b_{3}}^{-1}\in G_{2},\ \ \mathrm{and}$
@@
Moreover, by (9), (23) and (26), we obtain the following:
(27) $t_{a_{1}}^{-1}t_{b_{1}}=t_{a_{3}}t_{a_{1}}^{-1}\cdot(t_{a_{3}}t_{b_{1}}^{-1})^{-1}\in G_{2},\ \ \mathrm{and}$
(28) $t_{a_{2}}t_{b_{1}}^{-1}=t_{a_{2}}t_{a_{3}}^{-1}\cdot t_{a_{3}}t_{b_{1}}^{-1}\in G_{2}.$
Here, when we set $\sigma_{n}=\rho_{n}\cdot t_{a_{1}}^{-1}t_{b_{1}}(t_{a_{2}}t_{b_{1}}^{-1})^{-n+1}$, which is in $G_{2}$ by (27) and (28), we see that $\sigma_{n}=rt_{c_{1}}t_{a_{2}}^{-n+1}$ by the definition of $\rho_{n}$. Since it is easy to check that $\phi^{-1}(a_{2},b_{1})=(a_{2},c_{3})$, the element $t_{a_{2}}t_{c_{3}}^{-1}$ is in $G_{2}$ by (28) and (11). Moreover, by $\sigma_{n}^{2}(a_{2},c_{3})=\sigma_{n}(a_{3},c_{1})=(a_{1},c_{2})$, the follwing holds:
$t_{a_{3}}t_{c_{1}}^{-1}\in G_{2},\ \ \mathrm{and}$